# Unveiling the optimization process of Physics Informed Neural Networks: How accurate and competitive can PINNs be?

### v1

@@ 3.1.1 Impact of the optimization algorithm
In a similar light, we can investigate the effect of using a slightly modified version of the usual MSE loss function (2). In particular, we explore the consequences of a general approach, evaluating the loss function through a user-defined monotonically increasing function $g$:
$\mathcal{J}_{g}=g\left(\mathcal{J}\right).$ (25)
Near the minimum, the second term in brackets can be neglected, because $\nabla\mathcal{J}\simeq 0$, and both Hessian matrices are proportionally related to each other with ratio $g^{\prime}(\mathcal{J})$. Two obvious choices for $\mathcal{J}_{g}$ are
$\mathcal{J}_{1/2}$ $\equiv\sqrt{\mathcal{J}},$ (27)
$\mathcal{J}_{\log}$ $\equiv\log{\mathcal{J}}.$ (28)
@@
Since both the square root and the natural logarithm exhibit derivatives exceeding 1 when $\mathcal{J}\ll 1$, they have the potential to accelerate convergence. We demonstrate this by training three identical networks with different loss functions, namely $\mathcal{J},\mathcal{J}_{1/2},\mathcal{J}_{\log}$. We use the standard BFGS algorithm in order to isolate the impact of the loss function modifications from that of the BFGS modifications.
@@
When $\mathcal{J}>1$ the loss function modifications could have the opposite effect, i.e. decelerate convergence because of the decreased slope of $\mathcal{J}_{g}$. To prevent this, we utilize $\mathcal{J}_{g}$ exclusively during the BFGS phase of training, while maintaining the use of $\mathcal{J}$ during the Adam stage. Figure 4 illustrates the effect of the loss function modifications on convergence. Please observe that, for comparison purposes, we represent the standard MSE loss $\mathcal{J}$ in the plot, while the actual training has been conducted using the corresponding $\mathcal{J}_{g}$. We achieve an improvement of roughly two orders of magnitude by evaluating the loss through a monotonic function $g$, everything else being equal. $\mathcal{J}_{\log}$ converges slightly faster than $\mathcal{J}_{1/2}$ because its slope is steeper close to the minimum. This enhancement becomes more evident when examining the relative error norms of the PDE solution $\mathcal{P}$ and its derivatives, as presented in table 5.
Figure 4: Loss function vs. iterations for the current-free Grad-Shafranov equation. Results are obtained for the loss function modifications considered in section 3.1.2. Note that we plot the MSE loss (2) and not the modifications (25) used during training. In all cases, the standard BFGS algorithm is used.
@@
As in the previous section, one can attribute the improved convergence to the better conditioning of the Hessian matrix. We employ the same optimization algorithm in all cases. However, since we minimize different loss functions, the inverse Hessian approximations vary at each iteration. Consequently, the preconditioned Hessian $\mathrm{hess}(\mathcal{J}_{g}(\bm{z}))$ may exhibit different condition numbers $\kappa$ depending on the selection of $g$. Indeed, by analyzing the eigenvalue spectra of figure 5, it is apparent that $\mathcal{J}_{\log}$ exhibits a smaller condition number compared to $\mathcal{J}$, with more eigenvalues clustered around unity and the majority of its eigenvalues concentrated at higher values. All these features suggest a better-conditioned Hessian matrix and explain its superior performance. Similar observations hold for $\mathcal{J}_{1/2}$.
Figure 5: Eigenvalue spectra of the preconditioned Hessian matrix $\mathrm{hess}\left(\mathcal{J}\left(\bm{z}\right)\right)$ for $\mathcal{J}$ (5(a)), $\mathcal{J}_{1/2}$ (5(b)), and $\mathcal{J}_{\mathrm{log}}$ (5(c)).
@@
Changes to the loss function can be implemented combined with the selection of a better optimization algorithm. For the range of choices that we have explored, the latter seem to have a more pronounced effect on convergence. Using a combination of the best choices in each case can lead to overall better results, as seen in figure 6. In practice, however, it is much simpler to change the loss function than to modify an existing optimizer or to develop a new one.
Figure 6: Loss function vs. iterations for the current-free Grad-Shafranov equation considering various combinations of optimization algorithm/loss function modifications.
@@
3.2 Non-linear force-free solutions with higher order multipoles (NLGS)
@@
Up to now, our emphasis has been on a relatively simple, linear problem as to illustrate how enhancements in the optimization process can significantly improve the solution accuracy. Moving forward, we now show that our results generalize nicely to more complex solutions of the Grad-Shafranov equation.
We consider a non-linear force-free magnetospheric model by introducing a toroidal function similar to the one chosen in [26] but generalized for negative values of $\mathcal{P}$. This is:
@@
where $(s,\mathcal{P}_{c},\sigma)$ are parameters that control the relative strength of the toroidal and the poloidal magnetic fields, the extent of the current-filled region, and the degree of non-linearity of the model, respectively. At the surface of the star, the boundary condition consists of eight multipoles, so that $b_{l\leq 8}\neq 0$ in equation (15).
@@
Figure 7: Magnetic field lines obtained for the force-free case (black). The field lines corresponding to the current-free case are also plotted for comparison.
Analytical solutions are not available for the non-linear Grad-Shafranov equation, thus making it impossible to calculate the error of our solution through direct comparison. Instead, we adopt the approach that we introduced in [16]: we create a regular grid $\left\{q_{i},\mu_{j}\right\}_{i,j=1}^{N_{0}}$, and we evaluate the PINN solution $\mathcal{P}$ at all points in the grid using a forward pass. Then, we discretize the PDE (13) using a second-order finite differences scheme. If $\mathcal{P}$ was the exact solution of the PDE, the $L_{2}$-error of the discretized PDE $\epsilon_{\mathrm{FD}}$ would decrease with increasing resolution as $\sim N_{0}^{-2}$. In reality, $\mathcal{P}$ is only an approximate solution with an intrinsic error $\epsilon_{\mathrm{NN}}$ inherited from the quality and accuracy of the PINN. Therefore, $\epsilon_{\mathrm{FD}}$ will follow this power law only up to the point where the PINN approximation error $\epsilon_{\mathrm{NN}}$ starts to be important compared to the discretization error $\epsilon_{\mathrm{FD}}$ and, as a consequence, it will stagnate to the value of $\epsilon_{\mathrm{NN}}$. If the PINN approximation is not accurate, this will happen at a relatively low resolution. On the other hand, if $\epsilon_{\mathrm{FD}}$ dominates $\epsilon_{\mathrm{NN}}$ for high resolutions, then the PINN should be very close to the exact solution.
Figure 8 shows the performance of the BFGS modifications (top panels) and of the loss function modifications (bottom panels) on this problem. The left panels depict the evolution of the loss function with the number of iterations, whereas the right panels illustrate the estimation of the real errors, as outlined previously ($L_{2}$-error of the discretized PDE as a function of the number of grid points $N_{0}$). We see that the improvement in convergence and precision is massive for this complicated problem just as for the simple problem of section 3.1.
@@
Figure 8: Left: Loss function vs. iterations for the non-linear Grad-Shafranov equation. Results are obtained considering the BFGS algorithm modifications with the standard loss function (upper) and for the loss function modifications with the standard BFGS algorithm (lower). Right: $L_{2}$ norm of the discretized PDE as a function of the grid resolution $N_{0}$ obtained for each case.
@@
The PINN error when the standard BFGS algorithm is used for training is about $\epsilon_{\mathrm{NN}}\sim 10^{-3}$. SSBFGS and SSBroyden achieve much lower errors. Specifically, $\epsilon_{\mathrm{NN}}$ begins to surpass $\epsilon_{\mathrm{FD}}$ at an approximate resolution of $4500\times 4500$ (not depicted in the figure), near the memory limit of the machine employed for generating the results in this paper. Note that, since our discretization scheme is second order, the absolute error of the PDE is $\epsilon_{\mathrm{NN}}^{2}$. In table 6 we gather detailed error metrics for all the BFGS algorithm / loss function modifications we have used. In brief, aiming for a comparable precision level ($<10^{-4}$) with a second-order finite difference scheme would require a grid comprising at least 1000x1000 points for this 2D problem. In higher dimensions, the advantages of PINNs would become even more pronounced.
@@ 3.3 Parameter study
Figure 9: Approximation error $\epsilon_{\mathrm{NN}}$ for different depths as a function of the neurons in every layer. Left panel (9(a)): comparison between using BFGS and $\mathcal{J}$ and SSBroyden with $\mathcal{J}$. Right panel (9(b)): comparison between using BFGS with $\mathcal{J}_{\mathrm{log}}$ and SSBroyden with $\mathcal{J}$.
@@ 4 Other physics problems
KdV 3 30 20 10 10 $\left[0,5\right]\times[0,20]$
@@
Table 2: Same as Table 2 for the physical applications in this section.
In figure 11 we show the evolution of the loss function with the number of iterations for all problems, reaffirming that the findings outlined in section 3 hold universally. In all cases, we use the Adam optimizer and the standard MSE loss $\mathcal{J}$ for the initial training phase (10000 iterations). After that, we train using either BFGS with $\mathcal{J}_{\log}$ or using SSBroyden with $\mathcal{J}$, which were the most competitive combinations of modifications in section 3. In C, we provide additional tables detailing the errors and precision attained in each scenario.
@@
In [27] they consider the simplest case with $k=1$. We decided to raise this number to $k=4$, in order to get a more pronounced oscillatory behavior, that would challenge our solver. Results for the loss function and the error estimations can be found in figure 11(a) and in table 7.
@@
$\frac{\partial v}{\partial t}-\frac{\partial^{2}u}{\partial x^{2}}-\left(u^{2}+v^{2}\right)u=0,$ (38)
$\frac{\partial u}{\partial t}+\frac{\partial^{2}v}{\partial x^{2}}+\left(u^{2}+v^{2}\right)v=0.$ (39)
@@
The loss function is defined as the sum of two terms, which we construct according to equation (2) by identifying $\mathcal{L}_{v}=\frac{\partial}{\partial t}-uv$, $G_{v}=\frac{\partial^{2}u}{\partial x^{2}}+u^{3}$ and $\mathcal{L}_{u}=\frac{\partial}{\partial t}+uv$, $G_{u}=-\frac{\partial^{2}v}{\partial x^{2}}-v^{3}$.
@@
Results for the loss function and the error estimations can be found in figure 11(b) and in table 8
@@
The Neumann condition (47) is imposed via soft-enforcement as an additional term in the loss function. The total loss function is therefore calculated as $\mathcal{J}=\mathcal{J}_{\mathrm{PDE}}+\frac{\lambda}{N_{b}}\sum_{i=1}^{N_{b}}\lVert\frac{\partial}{\partial_{x}}u(t,x_{0}+L)-g_{3}(t)\rVert^{2}$, where $N_{b}$ is the number of points considered at the boundary and $\lambda$ is a hyperparameter to balance both terms. We set $N_{b}=1000$ and $\lambda=5$, as we found accurate results with these particular choices.
@@
We choose an initial condition such that the two solitons have initial positions $x_{1}=-2,x_{2}=2$ and initial velocities $c_{1}=6,c_{2}=2$. Since $x_{1}<x_{2}$ but $c_{1}>c_{2}$, eventually, the solitons will collide, triggering a non-linear interaction between them. We selected the aforementioned values to ensure that this interaction is significant and observable within the time domain under consideration. Before or after this interaction, the solitons will travel as a linear superposition of waves (that is, as single solitons) with their respective velocities. Fig. 10 shows four snapshots of the solution to appreciate this evolution. Results for the loss function and the error estimations can be found in figure 11(c) and in table 9.
Figure 10: Snapshots of the two-soliton solution (43). Solid black line: PINN approximation. Red dots: analytical solution.

### v2

@@ 4.1.1 Impact of the optimization algorithm
In a similar light, we can examine the impact of utilizing the modified versions of the loss function function discussed in section 3.3. To this end, we train three identical networks with different loss functions, namely $\mathcal{J},\mathcal{J}_{1/2},\mathcal{J}_{\log}$. We use the standard BFGS algorithm in order to isolate the impact of the loss function modifications from that of the BFGS modifications.
@@
When $\mathcal{J}>1$, the loss function modifications could decelerate convergence because of the decreased slope of $\mathcal{J}_{g}$. To prevent this, we utilize $\mathcal{J}_{g}$ exclusively during the BFGS phase of training, while maintaining the use of $\mathcal{J}$ during the Adam stage. The right panel of figure 1 illustrates the effect of the loss function modifications on convergence. Please observe that, for comparison purposes, we represent the standard MSE loss $\mathcal{J}$ in the plot, while the actual training has been conducted using the corresponding $\mathcal{J}_{g}$. We achieve an improvement of roughly two orders of magnitude by evaluating the loss through a monotonic function $g$, everything else being equal. $\mathcal{J}_{\log}$ converges slightly faster than $\mathcal{J}_{1/2}$ because its slope is steeper close to the minimum. This enhancement becomes more evident when examining the relative error norms of the PDE solution $\mathcal{P}$ and its derivatives, as presented in table 2.
@@
As in the previous section, one can attribute the improved convergence to the better conditioning of the Hessian matrix. We employ the same optimization algorithm in all cases. However, since we minimize different loss functions, the inverse Hessian approximations vary at each iteration. Consequently, the preconditioned Hessian $\mathrm{hess}(\mathcal{J}_{g}(\bm{z}))$ may exhibit different condition numbers $\kappa$ depending on the selection of $g$. Indeed, by analyzing the eigenvalue spectra in the middle and lower right panels of figure 2, it is apparent that $\mathcal{J}_{\log}$ exhibits a smaller condition number compared to $\mathcal{J}$, with the majority of its eigenvalues concentrated at higher values and more eigenvalues clustered around unity. All these features suggest a better-conditioned Hessian matrix and explain its superior performance. Similar observations hold for $\mathcal{J}_{1/2}$.
@@
Changes to the loss function can be implemented combined with the selection of a better optimization algorithm. For the range of choices that we have explored, the latter seem to have a more pronounced effect on convergence. Using a combination of the best choices in each case can lead to overall better results. In practice, however, it is much simpler to change the loss function than to modify an existing optimizer or to develop a new one.
Table 2: Relative error (equation (89)) of the flux function $\mathcal{P}$ and the magnetic field components $B_{r},B_{\theta}$ for the current-free Grad-Shafranov equation, employing different combinations of optimizer/loss function. The errors are averaged over multiple training runs. These runs employ different random initializations of the trainable parameters to ensure robustness. The results are presented in the format (mean ± standard deviation), giving a clear indication of both the average error and the variability across the different trials.
Optimizer Loss $E^{(2)}_{\mathcal{P}}$ $E^{(2)}_{B_{r}}$ $E^{(2)}_{B_{\theta}}$
BFGS $\mathcal{J}$ $(2.8\pm 1.3)\times 10^{-6}$ $(2.1\pm 1.2)\times 10^{-5}$ $(9\pm 2)\times 10^{-6}$
BFGS $\mathcal{J}_{1/2}$ $(3.6\pm 0.9)\times 10^{-7}$ $(4\pm 2)\times 10^{-6}$ $(1.1\pm 0.3)\times 10^{-6}$
BFGS $\mathcal{J}_{\mathrm{log}}$ $(2.3\pm 0.9)\times 10^{-7}$ $(2.4\pm 0.9)\times 10^{-6}$ $(7\pm 3)\times 10^{-7}$
SSBFGS $\mathcal{J}$ $(1.9\pm 0.4)\times 10^{-7}$ $(2.3\pm 0.7)\times 10^{-6}$ $(4.5\pm 0.6)\times 10^{-7}$
SSBroyden $\mathcal{J}$ $(1.0\pm 0.2)\times 10^{-7}$ $(1.0\pm 0.3)\times 10^{-6}$ $(2.5\pm 0.7)\times 10^{-7}$
@@
4.2 Non-linear force-free solutions with higher order multipoles.
@@
Up to now, our emphasis has been on a relatively simple, linear problem as to illustrate how enhancements in the optimization process can significantly improve the solution accuracy. Moving forward, we now show that our results generalize nicely to more complex solutions, in particular, we consider the Non-Linear Grad-Shafranov (NLGS) equation. We introduce a toroidal function similar to the one chosen in [81] but generalized for negative values of $\mathcal{P}$. This is:
@@
where $(s,\mathcal{P}_{c},\sigma)$ are parameters that control the relative strength of the toroidal and the poloidal magnetic fields, the extent of the current-filled region, and the degree of non-linearity of the model, respectively. At the surface of the star, the boundary condition consists of eight multipoles, so that $b_{l\leq 8}\neq 0$ in equation (30). An example solution is shown in figure 3.
@@
Figure 3: Magnetic field lines obtained for the force-free case (black). The field lines corresponding to the current-free case (red) are also plotted for comparison.
Figure 4 shows the performance of the BFGS modifications (upper panels) and of the loss function modifications (lower panels) on this problem. The left panels display the evolution of the loss function as the number of iterations increases, while the right panels present the $L_{2}$-error of the discretized PDE as a function of the number of grid points in the test grid. The minimum of this error is used to estimate the overall error $\epsilon_{\mathrm{NN}}$, as detailed in C. The improvement in convergence and precision is substantial for this more complex problem, similar to the simpler problem of section 4.1. Table 3 shows the specific values of $\epsilon_{\mathrm{NN}}$ for the different choices of optimizer and loss function.
@@
Figure 4: Evolution of the loss function with training iterations (left panels) and $L_{2}$ norm of the discretized PDE (see C) as a function of the grid resolution $N_{0}$ (right panels) for the non-linear Grad-Shafranov equation. The upper panels correspond to the optimizer modifications while the lower panels correspond to the loss function modifications.
@@
The PINN error when the standard BFGS algorithm is used for training is about $\epsilon_{\mathrm{NN}}\sim 10^{-3}$. SSBFGS and SSBroyden achieve much lower errors, of the order of $10^{-5}$. Specifically, $\epsilon_{\mathrm{NN}}$ begins to surpass $\epsilon_{\mathrm{FD}}$ at an approximate resolution of $4500\times 4500$ (not depicted in the figure), near the memory limit of the machine employed for generating the results in this paper. Note that, since our discretization scheme is second order, the absolute error of the PDE is $\epsilon_{\mathrm{NN}}^{2}$. In summary, aiming for a comparable precision level ($<10^{-4}$) with a second-order finite difference scheme would require a grid comprising at least $1000\times 1000$ points for this 2D problem. In higher dimensions, the advantages of PINNs would become even more pronounced.
Table 3: PINN approximation error $\epsilon_{\mathrm{NN}}$ (equation (91)) for the non-linear Grad-Shafranov equation. The errors are averaged over multiple training runs. These runs employ different random initializations of the trainable parameters to ensure robustness. The results are presented in the format (mean ± standard deviation), giving a clear indication of both the average error and the variability across the different trials.
Optimizer Loss $\epsilon_{\mathrm{NN}}$
BFGS $\mathcal{J}$ $(1.8\pm 0.2)\times 10^{-3}$
BFGS $\mathcal{J}_{1/2}$ $(4.0\pm 0.7)\times 10^{-4}$
BFGS $\mathcal{J}_{\mathrm{log}}$ $(7\pm 4)\times 10^{-5}$
SSBFGS $\mathcal{J}$ $(4.0\pm 1.2)\times 10^{-5}$
SSBroyden $\mathcal{J}$ $(1.5\pm 0.2)\times 10^{-5}$
@@ 4.3 Parameter study
Figure 5: Approximation error $\epsilon_{\mathrm{NN}}$ for different depths as a function of the neurons in each layer. Left panel: comparison between using BFGS and $\mathcal{J}$ and SSBroyden with $\mathcal{J}$. Right panel: comparison between using BFGS with $\mathcal{J}$ and BFGS with $\mathcal{J}_{\mathrm{log}}$
@@ 5 Other physics problems
Table 4: Architecture and training hyperparameters for various physical applications consdidered in section 5. In all cases, we use a $\tanh$ activation function for the hidden layers. The Neurons column refers to neurons per hidden layer. The Adam it. column refers to the number of iterations where the Adam optimizer is used before switching to a quasi-Newton method. The Batch size column refers to the number of points sampling the domain for each training set. The training set changes every 500 iterations in order to sample as many points as possible. The Iterations column denote the number of Quasi-Newton iterations.
@@
KdV 3 30 20 10 15 $\left[0,5\right]\times[0,20]$
1DB 3 20 10 5 10 $\left[0,1\right]\times\left[-1,1\right]$
AC 3 30 20 5 10 $\left[0,1\right]\times\left[-1,1\right]$
@@
LDC 6 20 20 0 25 $[0,1]\times[0,1]$
In all cases, we use the Adam optimizer and the standard MSE loss $\mathcal{J}$ for the initial training phase. After that, we train using either BFGS with $\mathcal{J}_{\log}$ or using SSBroyden with $\mathcal{J}$, which were the most competitive combinations of modifications in section 4.
For each problem, we show the evolution of the loss function with the number of iterations and the evolution of the relative $L_{2}$ error with respect to some reference solution, reaffirming that the findings outlined in section 4 hold universally. In addition, we present a comparison of the PINN prediction against the reference solution (the exact solution when available) and the distribution of errors in the domain.
2D Helmholtz equation (2DH)
We begin by examining the 2D Helmholtz equation problem as outlined in [57, 82]:
$\nabla^{2}u+k^{2}u-q(x,y)=0,$ (37)
where $k$ is a constant, and the source term
$q(x,y)=-\sin(\pi a_{1}x)\sin(\pi a_{2}y)\left[\pi^{2}\left(a_{1}^{2}+a_{2}^{2}\right)-k^{2}\right],$ (38)
is chosen such as the solution to the problem is analytical
$u(x,y)=\sin\left(\pi a_{1}x\right)\sin\left(\pi a_{2}y\right),$
where $a_{1},a_{2}\in\mathbb{Z}$. The computational domain is the square $\left[-1,1\right]\times\left[-1,1\right]\in\mathbb{R}^{2}$.
Imposing periodic boundary conditions in the $x$ and $y$ directions, the PINN solution is
$u(x,y)=\mathcal{N}\left(\cos\pi x,\sin\pi x,\cos\pi y,\sin\pi y\right),$ (39)
where $\mathcal{N}$ is the output of the network. The loss function is then constructed using equation (2), where $\mathcal{L}=\nabla^{2}$ and $G=-k^{2}u+q(x,y)$. We consider a low wavenumber case with $a_{1}=1$, $a_{2}=4$ and $k=1$, as in [57]. We remark that, if we consider periodic boundary conditions, the constant $k$ should be chosen such that the solution of the homogeneous PDE with periodic boundary conditions is zero. It can be readily seen that the homogeneus solution is zero, provided that $k^{2}\neq\pi^{2}\left(n^{2}+m^{2}\right)$, where $n,m\in\mathbb{Z}$.
We refer to table 4 for the specific set of hyperparameters chosen for this problem. The training set follows a random uniform distribution in the domain and is resampled every $500$ iterations. Figure 6 shows the evolution of the loss function with iterations, employing the BFGS algorithm in conjunction with the MSE loss, the SSBroyden algorithm with the MSE loss, and the BFGS algorithm with the logarithm of the MSE loss $\mathcal{J}_{\mathrm{\log}}$. We notice again that the modifications of the BFGS algorithm and the loss function introduce a remarkable improvement in convergence, achieving a reduction of $\sim 2-3$ orders of magnitude at the end of the training process. This is reflected in the relative $L_{2}$ error, where our modifications achieved to reduce the error between $1$ and $2$ orders of magnitude. Figure 7 shows the analytical and the PINN solutions, together with the absolute difference between them.
Figure 6: Convergence plots for the 2D Helmholtz equation with low wavenumber ($a_{1}=1$, $a_{2}=4$). Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error with respect to the reference solution on a fixed grid.
Figure 7: Solution comparison for the 2D Helmholtz equation with low wavenumber ($a_{1}=1$, $a_{2}=4$). Left panel: reference solution. Middle panel: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panel: absolute difference between the two. Note that the scale of the absolute error is $10^{-8}$.
We have also considered a more challenging case with higher wavenumber solution, where $a_{1}=a_{2}=6$ and $k=1$ as in [82]. As expected, resolving smaller structures calls for increased network complexity (see table 4) to achieve accurate results, but the same improvements of the previous cases are observed in figure 8. Once again, our modifications achieved to reduce the error significantly, obtaining a $L_{2}$ relative error of roughly one order of magnitude. In figure 9 we show the analytical and the PINN solutions (employing the self-scaled Broyden algorithm), and the relative difference between them. Errors are summarized later in table 5.
Figure 8: Convergence plots for the 2D Helmholtz equation with high wavenumber ($a_{1}=a_{2}=6$). Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error with respect to the reference solution on a fixed grid.
Figure 9: Solution comparison for the 2D Helmholtz equation with high wavenumber ($a_{1}=a_{2}=6$). Left panel: reference solution. Middle panel: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panel: absolute difference between the two.
@@
In [83] they consider the simplest case with $k=1$. We decided to raise this number to $k=4$, in order to get a more pronounced oscillatory behavior, that would challenge our solver.
Results for the loss function and the error norms are shown in figure 10 and table 5 respectively. Figure 11 shows the analytical solution, the PINN solution and their absolute difference, obtained with the self-scaled Broyden algorithm and the standard MSE loss.
Figure 10: Convergence plots for the non-linear Poisson equation. Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error with respect to the reference solution on a fixed grid.
Figure 11: Solution comparison for the non-linear Poisson equation. Left panel: reference solution. Middle panel: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panel: absolute difference between the two.
@@
$\frac{\partial v}{\partial t}-\frac{1}{2}\frac{\partial^{2}u}{\partial x^{2}}-\left(u^{2}+v^{2}\right)u=0,$ (48)
$\frac{\partial u}{\partial t}+\frac{1}{2}\frac{\partial^{2}v}{\partial x^{2}}+\left(u^{2}+v^{2}\right)v=0.$ (49)
@@
The loss function is defined as the sum of two terms, which we construct according to equation (2) by identifying $\mathcal{L}_{v}=\frac{\partial}{\partial t}-uv$, $G_{v}=\frac{1}{2}\frac{\partial^{2}u}{\partial x^{2}}+u^{3}$ and $\mathcal{L}_{u}=\frac{\partial}{\partial t}+uv$, $G_{u}=-\frac{1}{2}\frac{\partial^{2}v}{\partial x^{2}}-v^{3}$.
@@
Results for the loss function can be found in figure 12. In this case, while no analytical solution exists, the periodicity of the boundary conditions enables the computation of a highly accurate numerical solution using spectral methods. Specifically, we employed the Chebfun package in MATLAB (see [95] for details) with a spectral Fourier discretization of 3000 modes in combination with the EDTRK4 algorithm for the temporal part (see [96]), choosing a step size of $\frac{\pi}{2}10^{-6}$. The relative $L_{2}$ norms between the PINN and the reference solution can be found in table 5. Figure 13 shows the numerical solution of $\left|\Psi\right|$, the PINN solution calculated employing the self-scaled Broyden algorithm, and the absolute difference between the two solutions, which are of the order of $10^{-5}$.
Figure 12: Convergence plots for the non-linear Schrödinger equation. Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error for $\left|\Psi\right|$ with respect to the reference solution on a fixed grid.
Figure 13: Solution comparison for the non-linear Schrödinger equation. The colormaps correspond to the module of the wave function. Left panel: reference solution. Middle panel: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panel: absolute difference between the two. The solution is shown between $\left[-5,5\right]$, as outside this interval the solution decays rapidly to zero.
@@
The Neumann condition (57) is imposed via soft-enforcement as an additional term in the loss function. The total loss function is therefore calculated as $\mathcal{J}=\mathcal{J}_{\mathrm{PDE}}+\frac{\lambda}{N_{b}}\sum_{i=1}^{N_{b}}|\frac{\partial}{\partial_{x}}u(t,x_{0}+L)-g_{3}(t)|^{2}$, where $N_{b}$ is the number of points considered at the boundary and $\lambda$ is a hyperparameter to balance both terms. We set $N_{b}=1000$ and $\lambda=5$, as we found accurate results with these particular choices.
@@
We choose an initial condition such that the two solitons have initial positions $x_{1}=-2,x_{2}=2$ and initial velocities $c_{1}=6,c_{2}=2$. Since $x_{1}<x_{2}$ but $c_{1}>c_{2}$, eventually, the solitons will collide, triggering a non-linear interaction between them. We selected the aforementioned values to ensure that this interaction is significant and observable within the time domain under consideration. Before or after this interaction, the solitons will travel as a linear superposition of waves (that is, as single solitons) with their respective velocities.
Figure 14: Convergence plots for the Korteweg-De Vries equation. Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error with respect to the reference solution on a fixed grid.
Figure 15: Solution comparison for the Korteweg-De Vries equation. Left panel: reference solution. Middle panel: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panel: absolute difference between the two.
Results for the loss function and the error norms are shown in figure 14 and table 5. Figure 15 shows the reference solution, the PINN solution obtained with the self-scaled Broyden algorithm, and the absolute difference between them.
One-dimensional Burgers equation (1DB)
A classical benchmark in the literature of PINNs ([1, 28, 82, 101]) is the viscous Burgers equation
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},$ (61)
with initial data $u_{0}(x)=-\sin{(\pi x)}$, homogeneous boundary conditions in the domain $\left(t,x\right)\in\left[0,1\right]\times\left[-1,1\right]$, and viscosity $\nu=0.01/\pi$. Here, the input of the network consists of the coordinates $(t,x)$ in the domain, whereas the initial and the boundary conditions are imposed by setting
$f_{b}(t,x)$ $=u_{0}(x)=-\sin\pi x,$ (62)
$h_{b}(t,x)$ $=t(x^{2}-1).$ (63)
The loss function is given by (2), where $\mathcal{L}=\frac{\partial}{\partial t}+u\frac{\partial}{\partial x}-\frac{\partial^{2}}{\partial x^{2}}$ and $G=0$. The specific set of hyperparameters for this problem can be found in table 4. The training set follows a random uniform distribution and is resampled every 500 iterations.
Figure 16 shows the evolution of the loss function and the relative $L_{2}$ error (left and right panels, respectively) for the three combinations of optimizer/loss function employed. The dashed horizontal line represents, to the best of our knowledge, the best result reported in the literature for this problem [102]. As in other recent studies, we compare the PINN solution against a numerical solution obtained using a spectral method. We once again utilized the Chebfun package, with a spectral Fourier discretization of 2000 modes combined with the EDTRK4 algorithm for the temporal part, choosing a step size of $10^{-5}$. Final values for the relative $L_{2}$ errors are given in Table 5. Finally, figure 17 shows the reference and the PINN solutions, and the absolute difference between them, employing the self-scaled Broyden algorithm together with the MSE loss. The differences are at most of the order of $\sim 10^{-5}$ and particularly concentrated where the solution is very steep.
Figure 16: Convergence plots for the Burgers equation. Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error with respect to the reference solution on a fixed grid. The benchmark corresponds to results reported in [102].
Figure 17: Solution comparison for the Burgers equation. Left panel: reference solution. Middle panel: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panel: absolute difference between the two.
Allen-Cahn equation (AC)
Another important benchmark in PINN literature is the Allen-Cahn equation, which is a non-linear reaction-diffusion equation that is employed to describe processes related to phase separation in binary or multi-component alloys. It is given by the following PDE:
$\frac{\partial u}{\partial t}-\epsilon\frac{\partial^{2}u}{\partial x^{2}}=-f(u),$ (64)
where $f(u)$ is a non-linear source and $\epsilon$ is constant. A popular candidate for this function is $f(u)=\kappa(u^{3}-u)$, where $\kappa$ is a constant. The higher the ratio $\kappa/\epsilon$, the more pronounced the nonlinear behavior of the solution, involving sharp transitions that make the problem difficult to solve numerically. We consider the values $\kappa=5$ and $\epsilon=10^{-4}$ for the constants and the initial condition $u_{0}(x)=x^{2}\cos\pi x$, $x\in[-1,1]$, as these are standard choices in many studies (see [1, 24, 57, 103, 102]). For benchmark purposes, the initial condition is soft-enforced, whereas the periodic boundary conditions are hard-enforced by considering the Fourier functions $\left\{\cos\frac{2\pi x}{L},\sin\frac{2\pi x}{L}\right\}$ with $L=2$, as inputs for the spatial part. It is worth noting that in some of the previously mentioned studies, the input layer includes significantly more Fourier modes than in the case presented here (see for example [57] or [102]). The total loss function is given by
$\mathcal{J}=\mathcal{J}_{\mathrm{PDE}}+\frac{\lambda}{N_{b}}\lVert u-u_{0}\rVert_{(t=0,x)}^{2},$ (65)
with $\lambda=100$. The training process consists of $5000$ Adam iterations, followed by quasi-Newton training process with improved optimizers. The rest of hyperparameters are given in table 4.
Figure 18 shows the evolution of the loss function and the $L_{2}$ relative error between the PINN solution and a reference numerical solution obtained with the Chebfun package with $512$ Fourier modes and the EDTRK4 algorithm with a step size of $10^{-5}$ for the temporal part (same as the one employed in [57, 102]). The self-scaled Broyden algorithm achieves a $L_{2}$ relative error of $2.2\times 10^{-6}$, which is approximately one order of magnitude lower than the best result that we found in the literature [102]. Figure 19 shows colormaps of the reference solution, the PINN solution, and their absolute difference.
Figure 18: Convergence plots for the Allen-Cahn equation. Left panel: evolution of the loss function. Right panel: evolution of the relative $L_{2}$ error with respect to the reference solution on a fixed grid. The benchmark corresponds to results reported in [102].
Figure 19: Solution comparison for the Allen-Cahn equation. Left panel: reference solution. Middle panel: PINN prediction obtained with the SSBroyden algorithm and the standard MSE loss. Right panel: absolute difference between the two.